Determine the number of solutions and classify the type of solutions for each of the following equations. Justify your answer.
c) ~2x^2 – 4x - 7 ~2 = 0
(~ is to stand for the square root sign, I had no other ideal how to put it on there)

Answer by J2R2R(94)

~2x^2 - 4x - 7~2 = 0


~ means square root; +/- means plus or minus


a=~2; b=-4; c=-7~2


Using the b^2 - 4ac test, we have b^2 - 4ac = 16 -- 4*~2*7~2 = 16 + 56 = 72


Therefore as this is greater than 0 we have two real different roots.


The roots are (-b+/-~(b^2-4ac))/2a
Giving (4+/-~72)/2~2
which simplifies to ~2+/-3


The two roots are ~2+3 and ~2-3